# RTGS: Enabling Real-Time Gaussian Splatting on Mobile Devices Using Efficiency-Guided Pruning and Foveated Rendering

### v2

@@ Abstract.
Point-Based Neural Rendering (PBNR), i.e., the 3D Gaussian Splatting-family algorithms, emerges as a promising class of rendering techniques, which are permeating all aspects of society, driven by a growing demand for real-time, photorealistic rendering in AR/VR and digital twins. Achieving real-time PBNR on mobile devices is challenging.
This paper proposes RTGS, a PBNR system that for the first time delivers real-time neural rendering on mobile devices while maintaining human visual quality. RTGS combines two techniques. First, we present an efficiency-aware pruning technique to optimize rendering speed. Second, we introduce a Foveated Rendering (FR) method for PBNR, leveraging humans’ low visual acuity in peripheral regions to relax rendering quality and improve rendering speed. Our system executes in real-time (above 100 FPS) on Nvidia Jetson Xavier board without sacrificing subjective visual quality, as confirmed by a user study. The code is open-sourced at https://github.com/horizon-research/Fov-3DGS.
@@ 1. Introduction
Rendering technologies are infiltrating every facet of society. For instance, rendering is critical to enabling digital twins (juarez2021digital, ) in emerging areas such as smart cities, digital healthcare, and telepresence. Reinvigorated interests in Augmented/Virtual Reality (AR/VR) further heighten the demand for real-time, photorealistic rendering.
Point-Based Neural Rendering (PBNR), i.e., the Gaussian Splatting-family algorithms (Kerbl2023GaussianSplatting, ; fan2023lightgaussian, ; lee2024compact, ; fang2024mini, ; niemeyer2024radsplat, ; girish2023eagles, ), emerges as a new class of rendering solutions, which revitalizes the classic point-based rendering techniques (gross2011point, ; levoy1985use, ; pfister2000surfels, ; zwicker2001surface, ) using modern neural rendering methods (mildenhall2021nerf, ). PBNR, like previous neural rendering algorithms such as Neural Radiance Fields (NeRF), offers photorealistic rendering by learning light-matter interactions from data, but is significantly faster by replacing the compute-intensive Multilayer Perceptrons (MLPs) in NeRF with lightweight point-based rasterization.
Nevertheless, PBNR is still far from real-time on mobile devices, rendering generally below 10 Frames-Per-Second (FPS) on the mobile Volta GPU (xaviersoc, ). This paper introduces RTGS, which, for the first time, delivers real-time PBNR on mobile devices while maintaining human visual quality. RTGS combines two key ingredients.
@@
Much of the recent efforts on optimizing PBNR models focus on pruning (fan2023lightgaussian, ; lee2024compact, ; fang2024mini, ; niemeyer2024radsplat, ; girish2023eagles, ), which, while reducing the model size, does not bring significant speedups. This is because existing pruning methods are single-minded in reducing the sheer number of points while being agnostic to the actual computational cost. We find that different points in a PBNR model contribute differently to the overall computation. Instead, we propose an efficiency-aware pruning method that directly optimizes for the rendering/inference speed (Sec. 3).
@@
RTGS also exploits characteristics of human vision to improve performance (Sec. 4). Human vision acuity is poor in the visual periphery (wandell1995foundations, ), an opportunity that has long been exploited: one can speed up rendering by gradually reducing the rendering quality as the pixel eccentricity increases (i.e., as pixels are positioned more at the visual periphery) with impunity, a technique known as Foveated Rendering (FR) (patney2016towards, ; guenter2012foveated, ).
@@
We evaluate our method using both subjective human studies and objective measurements of performance and quality. Across 12 participants, the subjective rendering quality of our method is statistically no worse than that of Mini-Splatting-D  (fang2024mini, ), the PBNR method that is the current best in rendering quality. Compared to five state-of-the-art PBNR methods, we outperform all of them in both objective rendering quality (by up to 0.4 dB in PSNR) and rendering speed (by up to 7.4$\times$ on a mobile Volta GPU).
@@ 2.1. Point-Based Neural Rendering
PBNR is a class of neural rendering techniques, exemplified by the 3D Gaussian Splatting (3DGS) algorithm (Kerbl2023GaussianSplatting, ) and its descendants (fan2023lightgaussian, ; lee2024compact, ; fang2024mini, ; niemeyer2024radsplat, ; girish2023eagles, ). Compared to previous neural rendering techniques, a.k.a., the NeRF-family algorithms (mildenhall2021nerf, ; barron2022mip, ; muller2022instant, ; chen2022tensorf, ; sun2022direct, ), PBNR is fundamentally more efficient (e.g., usually over 1,000 times faster), because it parameterizes the scene with discrete points (rather than voxels) to avoid redundant computations and renders via a lightweight rasterization-based process called splatting (rusinkiewicz2000qsplat, ; zwicker2001surface, ; ren2002object, ) rather than the heavy MLP inference.
@@
Projection.  Each ellipsoid is first projected/splatted to an ellipse on the image plane [^1] We use “points”, “ellipses”, and “ellipsoids” interchangeably, since there is a one-to-one mapping between them.. In the example of Fig. 1, the ellipsoids AE in the scene are splatted to ellipses ae on the image plane. The goal is to identify, for each pixel tile (e.g., $2\times 2$), which ellipses intersect with the tile and thus contribute to the pixel colors in the tile.
Sorting.  For each tile, we sort all the intersecting ellipses based on their depths to the image plane; that way, closer ellipses can be weighted more when calculating pixel colors. For instance in Fig. 1, ellipse c would be the closest.
Rasterization.  Finally, we calculate the intersections of all the ellipses in a tile with each pixel. The color of a pixel p is then computed using the classic volume rendering method (max1995volume_rendering, ; kaufman1993volume, ; levoy1988display, ), which integrates the contribution of all the intersecting ellipses from near to far:
@@
where $N$ is the number of ellipses intersecting p (i.e., a, c, d, e in Fig. 1), $\alpha_{i}$ is a function $f$ of various trainable parameters of the $i^{th}$ intersecting ellipse (e.g., opacity) and the camera pose, and $c_{i}$ is the color of the $i^{th}$ ellipse at the position of intersection, which is calculated using the Spherical Harmonics (SH) function $g$ with trainable coefficients $\text{SH}_{n}$. We refer readers to Kerbl et al. (Kerbl2023GaussianSplatting, ) for the details of $f$ and $g$.
@@ 2.2. Human Visual System and Foveated Rendering
It is well-known that human visual acuity drops as eccentricity increases, i.e., when objects are placed more toward the visual periphery (wandell1995foundations, ). This is due to a combination of larger pooling sizes (rodieck1985parasol, ; dacey1993mosaic, ) and a sparser photoreceptor distribution (curcio1990human, ; Song:2011:ConeDensity, ) on the retina as the eccentricity increases. Foveated Rendering (FR) (patney2016towards, ; guenter2012foveated, ) leverages this natural fall-off in visual acuity to speed up rendering by relaxing the rendering quality in peripheral regions. Fig. 2 illustrates an example where the visual content in high-eccentricity regions could be altered without being noticeable by users.
While in classic FR the peripherial rendering quality is relaxed by lowering the resolution, neural rendering offers another dimension: reducing the computational workload of each pixel. This new dimension is possible because the rendering load of each pixel is controlled by inferencing a deep learning model, which offers many knobs for accuracy-vs-speed trade-offs that have been extensively studied (han2015learning, ; blalock2020state, ; jin2020adabits, ; yang2019quantization, ). For instance, one can train a smaller model for rendering the visual periphery (deng2022fov, ). This paper will explore FR knobs unique to PBNR.
@@
A key question in FR is to determine how much quality to relax without introducing visual artifacts. It is well-established that commonly used visual quality metrics such as Peak Signal to Noise Ratio (PSNR) or Structural Similarity Index Measure (SSIM) (hore2010image, ) do not account for the eccentricity-dependent visual acuity drop in HVS (walton2021beyond, ; rosenholtz2016capabilities, ; strasburger2011peripheral, ) and, thus, are inadequate for FR: an image with a low PSNR at the visual periphery might not introduce visual artifacts. The altered image in Fig. 2, when placed in the visual periphery, is visually indiscriminable from the reference image.
This paper leverages an eccentricity-aware HVS Quality (HVSQ) metric (walton2021beyond, ) inspired by classic neuroscience studies about the human visual pathway (FreemanSimoncelli2011, ). Given a reference image, an altered image, and the eccentricity of each pixel (which depends on the display resolution and the eye-display distance), HVSQ quantifies how similar the two images are as viewed by humans; a lower HVSQ means more similar.
The principle behind the HVSQ metric is as follows. The retina aggregates photoreceptor outputs in spatial regions, called spatial poolings. In the image space, a spatial pooling corresponds to a set of adjacent pixels (e.g., SP in Fig. 2). The pooling size increases with eccentricity, usually quadratically. Computational models on HVS (walton2021beyond, ) show that as long as the statistics (mean and standard deviation) of the content in a spatial pooling between two images are close, humans can not discriminate between them. The statistics are calculated in a feature space (as opposed to the pixel space) to emulate the feature extraction in human’s early visual processing.
@@ 3.1. Motivations
The performance of recent PBNR models is far from real-time on mobile GPUs. Fig. 5 shows the FPS on the Mip-NeRF 360 (barron2022mip, ), Tanks&Temples (Knapitsch2017, ), and Deep Belending (hedman2018deep, ) dataset, measured on the mobile Volta GPU on Jetson Xavier across five recent PBNR models (fan2023lightgaussian, ; lee2024compact, ; fang2024mini, ; Kerbl2023GaussianSplatting, ). The data is plotted as a standard boxplot to show the FPS distribution across the 13 traces within the datasets.
3DGS (Kerbl2023GaussianSplatting, ) and Mini-Splatting-D (fang2024mini, ) are two dense models and generally are the slowest. Much of recent work focuses on pruning: reducing the number of points in a PBNR model (fan2023lightgaussian, ; lee2024compact, ; fang2024mini, ). While effective for reducing the model size, these methods do not significantly speed up rendering. The last three models in Fig. 5 are pruned models. While generally faster, they are still far below real-time, especially for immersive applications such as AR/VR, which would normally require an FPS of 75–90 (vive_pro2, ; meta_quest_pro, ; apple_vision_pro, ).
Why is Existing Pruning Ineffective?
Existing pruning methods focus on reducing the point count in a model, which is ineffective for improving speed in PBNR. To quantify this, Fig. 5 shows the inference latency ($x$-axis) vs. point count (left $y$-axis) of LightGS (fan2023lightgaussian, ) (which prunes 3DGS (Kerbl2023GaussianSplatting, )) trained on the bicycle trace in the Mip-NeRF 360 dataset at different pruning levels (between 75% and 97%). The latency reduction rate is slower than that of the point reduction rate.
@@ 4.1. Main Idea and Challenges
As with prior FR work (deng2022fov, ; guenter2012foveated, ; patney2016towards, ), we divide an image into $N$ regions (4 in the example), each corresponding to a quality level and is rendered by a separate model. The region currently under the user’s gaze has the highest quality ($𝑅1$ here). Lower-quality regions are rendered using lighter models, which are obtained by applying pruning and scale decay (Sec. 3) to a high-quality model.
@@
Blending is required in all FR algorithms (guenter2012foveated, ; patney2016towards, ). Due to the quality difference across levels, there is a sharp, undesirable boundary between two adjacent levels in the rendered image (a form of aliasing). To eliminate the boundary, a common technique is for each model to render slightly beyond its assigned boundary; thus, pixels at the boundary will be rendered twice and then are interpolated/blended to provide a smooth transition between the two levels.
@@
FR could increase the model size due to the need to store multiple models, exacerbating the storage pressure of PBNR models. For instance, the bicycle scene in the Mip-NeRF 360 dataset (barron2021mip, ) takes about 1.4 GB of space when trained with 3DGS (Kerbl2023GaussianSplatting, ); recent pruning methods  (fan2023lightgaussian, ) reduce the model size of that scene to about 490 MB, which is still large for mobile devices.

### v3

# MetaSapiens: Real-Time Neural Rendering with Efficiency-Aware Pruning and Accelerated Foveated Rendering

@@ Abstract.
Point-Based Neural Rendering (PBNR) is emerging as a promising class of rendering techniques, which are permeating all aspects of society, driven by a growing demand for real-time, photorealistic rendering in AR/VR and digital twins. Achieving real-time PBNR on mobile devices is challenging.
This paper proposes MetaSapiens, a PBNR system that for the first time delivers real-time neural rendering on mobile devices while maintaining human visual quality. MetaSapiens combines three techniques. First, we present an efficiency-aware pruning technique to optimize rendering speed. Second, we introduce a Foveated Rendering (FR) method for PBNR, leveraging humans’ low visual acuity in peripheral regions to relax rendering quality and improve rendering speed. Finally, we propose an accelerator design for FR, addressing the load imbalance issue in (FR-based) PBNR. Our evaluation shows that our system achieves an order of magnitude speedup over existing PBNR models without sacrificing subjective visual quality, as confirmed by a user study. The code and demo are available at: https://horizon-lab.org/metasapiens/.
Gaussian Splatting, Foveated Rendering, Hardware Accelerator
journalyear: 2025copyright: ccconference: Proceedings of the 29th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 1; March 30-April 3, 2025; Rotterdam, Netherlandsbooktitle: Proceedings of the 29th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 1 (ASPLOS ’25), March 30-April 3, 2025, Rotterdam, Netherlandsdoi: 10.1145/3669940.3707227isbn: 979-8-4007-0698-1/25/03ccs: Computer systems organization Architecturesccs: Human-centered computing Ubiquitous and mobile computing
@@ 1. Introduction
Rendering technologies are infiltrating every facet of society. For instance, rendering is critical to enabling digital twins (Juarez et al., 2021) in emerging areas such as smart cities, digital healthcare, and telepresence. Reinvigorated interests in Augmented/Virtual Reality (AR/VR) further heighten the demand for real-time, photorealistic rendering.
Point-Based Neural Rendering (PBNR), i.e., the Gaussian Splatting-family algorithms (Kerbl et al., 2023; Fan et al., 2024; Lee et al., 2024b; Fang and Wang, 2024; Niemeyer et al., 2024; Girish et al., 2024), is emerging as a new class of rendering solutions, which revitalizes the classic point-based rendering techniques (Gross and Pfister, 2011; Levoy and Whitted, 1985; Pfister et al., 2000; Zwicker et al., 2001) using modern neural rendering methods (Mildenhall et al., 2021). PBNR, like previous neural rendering algorithms such as Neural Radiance Fields (NeRF), offers photorealistic rendering by learning scene radiance from data, but is significantly faster by replacing the compute-intensive Multilayer Perceptrons (MLPs) in NeRF with lightweight point-based rasterization.
Nevertheless, PBNR is still far from real-time on mobile devices, rendering generally below 10 Frames-Per-Second (FPS) on the mobile Volta GPU (xav, 2018). This paper introduces MetaSapiens, which, for the first time, delivers real-time PBNR on mobile devices while maintaining human visual quality. MetaSapiens combines three key ingredients.
@@
Much of the recent efforts on optimizing PBNR models focus on pruning (Fan et al., 2024; Lee et al., 2024b; Fang and Wang, 2024; Niemeyer et al., 2024; Girish et al., 2024), which, while reducing the model size, does not bring significant speedups. This is because existing pruning methods are single-minded in reducing the sheer number of points while being agnostic to the actual computational cost. We find that different points in a PBNR model contribute differently to the overall computation. Instead, we propose an efficiency-aware pruning method that directly optimizes for the rendering/inference speed (Sec. 3).
@@
MetaSapiens also exploits characteristics of human vision to improve performance (Sec. 4). Human vision acuity is poor in the visual periphery (Wandell, 1995), an opportunity that has long been exploited: one can speed up rendering by gradually reducing the rendering quality as the pixel eccentricity increases (i.e., as pixels are positioned more at the visual periphery) with impunity, a technique known as Foveated Rendering (FR) (Patney et al., 2016; Guenter et al., 2012).
@@
Architectural Support.
While the two techniques above readily provide about an order of magnitude speedup on GPU, we co-design an accelerator architecture to further improve the performance (Sec. 5). Aside from providing hardware support for FR, the hardware addresses a key performance bottleneck in PBNR exacerbated by FR: low hardware utilization due to workload imbalance across tiles in a frame. Our hardware mitigates this issue via 1) a dynamic tile merging scheme that balances workloads across tiles and 2) incrementally pipelining adjacent stages through line-buffering.
@@
We evaluate our method using both subjective human studies and objective measurements of performance and quality. Across 12 participants, the subjective rendering quality of our method is statistically no-worse than that of Mini-Splatting-D  (Fang and Wang, 2024), a state-of-the-art PBNR method in rendering quality. Compared to five state-of-the-art PBNR methods, we out-perform all of them in both objective rendering quality (by up to 0.4 dB in PSNR) and rendering speed (by up to 7.4$\times$ on a mobile Volta GPU and 20.9$\times$ with our hardware support).
@@
  • We co-design an accelerator architecture, which addresses the load imbalance issue in PBNR and futher improves performance.
@@ 2.1. Point-Based Neural Rendering
PBNR is a class of neural rendering techniques, exemplified by the 3D Gaussian Splatting (3DGS) algorithm (Kerbl et al., 2023) and its descendants (Fan et al., 2024; Lee et al., 2024b; Fang and Wang, 2024; Niemeyer et al., 2024; Girish et al., 2024). Compared to previous neural rendering techniques, a.k.a., the NeRF-family algorithms (Mildenhall et al., 2021; Barron et al., 2022; Müller et al., 2022; Chen et al., 2022b; Sun et al., 2022), PBNR is fundamentally more efficient (e.g., usually over 1,000 times faster), because it parameterizes the scene with discrete points (rather than voxels) to avoid redundant computations and renders via a lightweight rasterization-based process called splatting (Rusinkiewicz and Levoy, 2000; Zwicker et al., 2001; Ren et al., 2002) rather than the heavy MLP inference.
@@
Projection.  Each ellipsoid is first projected/splatted to an ellipse on the image plane [^1] We use “points”, “ellipses”, and “ellipsoids” interchangeably: there is a one-to-one mapping between them.. In the example of Fig. 1, the ellipsoids AE in the scene are splatted to ellipses ae on the image plane. The goal is to identify, for each pixel tile (e.g., $2\times 2$), which ellipses intersect with the tile and thus contribute to the pixel colors in the tile.
Sorting.  For each tile, we sort all the intersecting ellipses based on their depths to the image plane; that way, closer ellipses can be integrated first when calculating pixel colors. For instance in Fig. 1, ellipse c would be the closest.
Rasterization.  Finally, we calculate the intersections of all the ellipses in a tile with each pixel. The color of a pixel p is then computed using the classic volume rendering method (Max, 1995; Kaufman et al., 1993; Levoy, 1988), which integrates the contribution of all the intersecting ellipses from near to far:
@@
where $N$ is the number of ellipses intersecting p (i.e., a, c, d, e in Fig. 1), $\alpha_{i}$ is a function $f$ of various trainable parameters of the $i^{th}$ intersecting ellipse (e.g., opacity) and the camera pose, and $c_{i}$ is the color of the $i^{th}$ ellipse at the position of intersection, which is calculated using the Spherical Harmonics (SH) function $g$ with trainable coefficients $\text{SH}_{n}$. We refer readers to Kerbl et al. (Kerbl et al., 2023) for the details of $f$ and $g$.
@@ 2.2. Human Visual System and Foveated Rendering
It is well-known that human visual acuity drops as eccentricity increases, i.e., when objects are placed more toward the visual periphery (Wandell, 1995). This is due to a combination of larger pooling sizes (Rodieck et al., 1985; Dacey, 1993) and a sparser photoreceptor distribution (Curcio et al., 1990; Song et al., 2011) on the retina as the eccentricity increases. Foveated Rendering (FR) (Patney et al., 2016; Guenter et al., 2012) leverages this natural fall-off in visual acuity to speed up rendering by relaxing the rendering quality in peripheral regions. Fig. 2 illustrates an example where the visual content in high-eccentricity regions could be altered without being noticeable by users.
While in classic FR the peripherial rendering quality is relaxed by lowering the resolution, neural rendering offers another dimension: reducing the computational workload of each pixel. This new dimension is possible because the rendering load of each pixel is controlled by inferencing a deep learning model, which offers many knobs for accuracy-vs-speed trade-offs that have been extensively studied (Han et al., 2015; Blalock et al., 2020; Jin et al., 2020; Yang et al., 2019). For instance, one can train a smaller model for rendering the visual periphery (Deng et al., 2022). This paper will explore FR knobs unique to PBNR.
@@
A key question in FR is to determine how much quality to relax without introducing visual artifacts. It is well-established that commonly used visual quality metrics such as Peak Signal to Noise Ratio (PSNR) or Structural Similarity Index Measure (SSIM) (Hore and Ziou, 2010) do not account for the eccentricity-dependent visual acuity drop in HVS (Walton et al., 2021; Rosenholtz, 2016; Strasburger et al., 2011) and, thus, are inadequate for FR: an image with a low PSNR at the visual periphery might not introduce visual artifacts. The altered image in Fig. 2, when placed in the visual periphery, is visually indiscriminable from the reference image.
This paper leverages an eccentricity-aware HVS Quality (HVSQ) metric (Walton et al., 2021) inspired by classic neuroscience studies about the human visual pathway (Freeman and Simoncelli, 2011). Given a reference image, an altered image, and the eccentricity of each pixel (which depends on the display resolution and the eye-display distance), HVSQ quantifies how similar the two images are as viewed by humans; a lower HVSQ means more similar.
The principle behind the HVSQ metric is as follows. The retina aggregates photoreceptor outputs in spatial regions, called spatial poolings. In the image space, a spatial pooling corresponds to a set of adjacent pixels (e.g., SP in Fig. 2). The pooling size increases with eccentricity, usually quadratically. Computational models on HVS (Walton et al., 2021) show that as long as the statistics (mean and standard deviation) of the content in a spatial pooling between two images are close, humans can not discriminate between them. The statistics are calculated in a feature space (as opposed to the pixel space) to emulate the feature extraction in human’s early visual processing.
@@ 3.1. Motivations
The performance of recent PBNR models is far from real-time on mobile GPUs. Fig. 5 shows the FPS on the Mip-NeRF 360 (Barron et al., 2022), Tanks&Temples (Knapitsch et al., 2017), and Deep Belending (Hedman et al., 2018) dataset, measured on the mobile Volta GPU on Jetson Xavier across five recent PBNR models (Fan et al., 2024; Lee et al., 2024b; Fang and Wang, 2024; Kerbl et al., 2023). The data is plotted as a standard boxplot to show the FPS distribution across the 13 traces within the datasets.
3DGS (Kerbl et al., 2023) and Mini-Splatting-D (Fang and Wang, 2024) are two dense models and generally are the slowest. Much of recent work focuses on pruning: reducing the number of points in a PBNR model (Fan et al., 2024; Lee et al., 2024b; Fang and Wang, 2024). While effective for reducing the model size, these methods do not significantly speed up rendering. For instance, CompactGS, LightGS, and Mini-Splatting in Fig. 5 are all pruned models; while generally faster than dense models, they are still below real-time, especially for immersive applications such as AR/VR, which would normally require an FPS of 75–90 (viv, 2021; met, 2022; app, 2024).
Why is Existing Pruning Insufficient?
Existing pruning methods focus on reducing the point count in a model, which is ineffective for improving speed in PBNR. To quantify this, Fig. 5 shows the inference latency ($x$-axis) vs. point count (left $y$-axis) of LightGS (Fan et al., 2024) (which prunes 3DGS (Kerbl et al., 2023)) trained on the bicycle trace in the Mip-NeRF 360 dataset at different pruning levels (between 75% and 97%). The latency reduction rate is slower than that of the point reduction rate.
@@ 4.1. Main Idea and Challenges
As with prior FR work (Deng et al., 2022; Guenter et al., 2012; Patney et al., 2016), we divide an image into $N$ regions (4 in the example), each corresponding to a quality level and is rendered by a separate model. The region currently under the user’s gaze has the highest quality ($𝑅1$ here). Lower-quality regions are rendered using lighter models, which are obtained by applying pruning and scale decay (Sec. 3) to a high-quality model.
@@
Blending is required in all FR algorithms (Guenter et al., 2012; Patney et al., 2016). Due to the quality difference across levels, there is a sharp, undesirable boundary between two adjacent levels in the rendered image (a form of aliasing). To eliminate the boundary, a common technique is for each model to render slightly beyond its assigned boundary; thus, pixels at the boundary will be rendered twice and then are interpolated/blended to provide a smooth transition between the two levels.
@@
FR could increase the model size due to the need to store multiple models, exacerbating the storage pressure of PBNR models. For instance, the bicycle scene in the Mip-NeRF 360 dataset (Barron et al., 2021) takes about 1.4 GB of space when trained with 3DGS (Kerbl et al., 2023); recent pruning methods  (Fan et al., 2024) reduce the model size of that scene to about 490 MB, which is still large for mobile devices.
@@ 4.2. Efficient Data Representation
Fig. 8. The overall architecture design. The the basic pipeline is similar to that of GSCore (Lee et al., 2024a), a recent PBNR accelerator. We augment the baseline to support FR (yellow-colored) and to address the workload imbalance issue in PBNR/FR (blue-colored).